OpenStudy (anonymous):

Solve the equation. √ ̅3x - 5 = 4

OpenStudy (anonymous):

\[\sqrt{x}-5=4\] or \[\sqrt{x-5}=4\]?

OpenStudy (anonymous):

first one solve via \[\sqrt{x}=9\] \[x=81\] second one solve via \[x-5=16\] \[x=21\]

OpenStudy (anonymous):

It's a long radical sign I'm sorry

OpenStudy (anonymous):

long radical sign 3x - 5 = 4